OpenStudy (anonymous):

Find the volume formed by rotating the region enclosed by: x = 10 y and x = y^3 with y≥0 about the y-axis

OpenStudy (callisto):

I guess it's similar to the previous one... 1. Put x=10y into x=y^3 => y(y^2 -10) =0 => y=0, y=-sqrt10 or y=sqrt10 2. \[Volume = 2\pi \int_{0}^{\sqrt{10}} y(y^3) -y(10y) dy \] Can you do it from here?

OpenStudy (anonymous):

I'm getting a negative value :s

OpenStudy (anonymous):

oops nevermind I forgot to put - sqrt10

OpenStudy (callisto):

It's weird.... Let me think about it... Sorry :(

OpenStudy (anonymous):

I don't think it's the right integral, I got 264.922 and its not working

OpenStudy (callisto):

Oh.... My fault :( \[volume = \pi \int_{0}^{\sqrt{10}}(y^3)^2 -(10y)^2dy\] Let see if it works this time :S

OpenStudy (callisto):

Something wrong with the limit...

OpenStudy (anonymous):

That one worked! using -sqrt 10 :-)

OpenStudy (callisto):

I see what's wrong. it should be \[volume = pi \int_{0}^{\sqrt{10}} (10y)^2 - (y^3)^2 dy\] This is the correct one :S

OpenStudy (callisto):

since x=10y has a longer distance to y-axis that x=y^3, so it should be (10y)^2 - (y^3)^2

OpenStudy (anonymous):

Sweet, thanks for your help!

OpenStudy (callisto):

Welcome :)